OpenStudy (anonymous):

What is the solution of 2(x - 5) + 4 = 30?

OpenStudy (anonymous):

You need to simplify the problem first. \[2(x-5) = 2x-10\] \[2x-10+4=30\] \[2x-10+4-4=30-4\] \[2x-10=26\] \[2x-10+10=26+10\] \[2x=36\] \[2x \div2=36\div2\] \[x=18\]

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

Yep!